Is (3, 2) a solution to this system of equations?
8x + 7y = 37
2y = 4x - 2

Answer by Alan3354(69443)

Is (3, 2) a solution to this system of equations?
8x + 7y = 37
2y = 4x - 2
-------------
Sub 3 for x and 2 for y and see.
8x + 7y = 37
8*3 + 7*2 =? 37
24 + 14 =? 37 not a solution.